a sample of crude oil has a densityof 0.87g/mL.what volume in liters does a 3.6 kg sample of this oil occupy?
1
Expert's answer
2012-07-05T11:22:34-0400
d = 0.87g/mL = 870g/L = 0.87kg/L
V = m / d = 3.6 / 0.87 = 4.1379L